# Is Algorithmic Stability Testable? A Unified Framework under Computational Constraints

### v1

@@ 1 Introduction
That is, $\mathcal{A}$ takes as input $\mathcal{D}=\{(X_{i},Y_{i})\}_{i\in[n]}$, a data set of any size $n$, and returns a fitted model ${\widehat{f}}_{n}:\mathcal{X}\rightarrow\mathbb{R}$, which maps a feature value $x$ to a fitted value, or prediction, ${\widehat{y}}={\widehat{f}}_{n}(x)$. Algorithmic stability is a notion that quantifies the sensitivity of an algorithm to small changes in the training data $\mathcal{D}$. Since its introduction (Devroye and Wagner,, 1979; Rogers and Wagner,, 1978; Kearns and Ron,, 1997), algorithmic stability has been a central notion in machine learning and statistics. Intuitively, it is a desirable property for learning algorithms—but it also leads to many downstream implications, as well. For instance, it is known that stable algorithms generalize (Kearns and Ron,, 1997; Bousquet and Elisseeff,, 2002; Poggio et al.,, 2004; Shalev-Shwartz et al.,, 2010). It also plays a critical role in a number of different problems, such as model selection (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013; Ren et al.,, 2023), reproducibility (Yu,, 2013), predictive inference (Steinberger and Leeb,, 2023; Barber et al.,, 2021; Liang and Barber,, 2023), selective inference (Zrnic and Jordan,, 2023) and inference for cross-validation error (Austern and Zhou,, 2020; Bayle et al.,, 2020; Kissel and Lei,, 2022).
@@ 1.1 Overview of contributions
In Section 2, we will present some background on algorithmic stability and black-box testing framework under computational constraints. Section 3 presents our main results on the limit of black-box tests for algorithmic stability. A brief discussion is provided in Section 4. We prove the main results in Section 5. Some technical details and extensions are deferred to the Appendix.
@@ 2 Background and problem setup
The fitted model ${\widehat{f}}_{n}=\mathcal{A}\big{(}\{(X_{i},Y_{i})\}_{i\in[n]};\xi\big{)}$ is now obtained by running algorithm $\mathcal{A}$ on data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, with randomization provided by the argument $\xi\in[0,1]$, which acts as a random seed. [^1] Throughout this paper, we implicitly assume measurability for all functions we define, as appropriate. For the algorithm $\mathcal{A}$, we assume measurability of the map $\left((x_{1},y_{1}),\ldots,(x_{n},y_{n}),\xi,x\right)\mapsto[\mathcal{A}(\{(x_{i},y_{i})\}_{i\in[n]};\xi)](x)$. Of course, the general definition (2) of a randomized algorithm also accommodates the deterministic (i.e., non-random) setting—we can simply choose an algorithm $\mathcal{A}$ whose output depends only on the data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, and ignores the second argument $\xi$. In this paper, we focus on the setting where $\mathcal{A}$ is symmetric in the training data—that is, permuting the order of the $n$ data points does not change the output of $\mathcal{A}$ (or the distribution of the output, if randomized—see Kim and Barber, (2023) for a formal definition).
@@ 2.1 Algorithmic stability and hypothesis testing formulation
There are many different notions of algorithmic stability in the literature (see, e.g., Bousquet and Elisseeff, (2002), Appendix A of Shalev-Shwartz et al., (2010) and Section 2.1.1 of Kim and Barber, (2023) for some comparisons). In this work, we follow Kim and Barber, (2023) and consider testing the following notion of stability:
@@ 2.2 Black-box tests with computational constraints
Compared to the definition of black-box test given in Kim and Barber, (2023), here we focus on black-box tests with computational constraints. A high-level illustration of a black-box test with computational constraints is provided in Figure 1; the formal definition is quite technical and we defer it to Appendix A. In particular, the computational budget for training is characterized by $B_{\textnormal{train}}$—the total number of data points that we are allowed to input to the algorithm $\mathcal{A}$, over all calls to $\mathcal{A}$ (for example, when running the test, we can choose to train $\mathcal{A}$ on a data set of size $n$, up to $B_{\textnormal{train}}/n$ many times). Some examples of black-box tests, and counterexamples (i.e., tests that do not follow the black-box framework), can be found in Kim and Barber, (2023, Section 2.3).
@@
Transparent models or black-box models?
Careful readers may notice that the version of black-box test we presented in Figure 1, is called a black-box test with transparent models by Kim and Barber, (2023). This means that while the training algorithm $\mathcal{A}$ can only be studied empirically via calls to $\mathcal{A}$ (that is, we cannot examine its theoretical properties, or, say, read the code that defines the implementation of $\mathcal{A}$), in contrast the fitted model ${\widehat{f}}^{(r)}$ returned by $\mathcal{A}$ at the $r$th round is transparent to us—we know its analytic expression. This is a less restrictive notion of black-box testing, giving more information to the analyst. In contrast, an even stronger constraint is the notion of a black-box test with black-box models in Kim and Barber, (2023), where the fitted models cannot be studied analytically and instead can only be studied via evaluating the models on finitely many test points.
In the main part of this paper, we will focus on the black-box test with transparent models, since this offers more flexibility to the analyst—and thus, establishing hardness of testing algorithmic stability in this setting is a stronger result. We extend our results to the black-box model setting in Appendix B.
3 Main results
In this section, we present the main results on a unified power upper bound of testing algorithmic stability for computationally constrained black-box tests. First, given any $(\mathcal{A},P,n)$, let us define
Theorem 1.
Fix any $\epsilon\geq 0$, $\delta\in[0,1)$. Let ${\widehat{T}}$ be any black-box test with computational budget $B_{\textnormal{train}}$ (as illustrated in Figure 1—see Appendix A for the formal definition). Suppose ${\widehat{T}}$ satisfies the assumption-free validity condition (3) at level $\alpha$.
Then for any $(\mathcal{A},P,n)$ that is $(\epsilon,\delta)$-stable (i.e., $\delta^{*}_{\epsilon}\leq\delta$), the power of ${\widehat{T}}$ is bounded as
$\mathbb{P}\{{\widehat{T}}=1\}\leq\alpha\cdot\min\Bigg{\{}\underbrace{\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{B_{\textnormal{train}}}{n}}}_{\begin{subarray}{c}\textnormal{computational constraint}\\ \textnormal{due to budget $B_{\textnormal{train}}$}\end{subarray}},\underbrace{\frac{\left(\frac{1-\delta^{*}_{\epsilon}}{1-\delta}\right)^{\frac{N_{\ell}}{n}}}{1-\frac{B_{\textnormal{train}}}{|\mathcal{Y}|}\wedge 1}}_{\begin{subarray}{c}\textnormal{statistical constraint due to}\\ \textnormal{limited samples in $\mathcal{Y}$}\end{subarray}},\underbrace{\frac{\left(\frac{1-\delta^{*}_{\epsilon}}{1-\widetilde{\delta}}\right)^{\frac{N_{\ell}+N_{u}}{n}}}{1-\frac{B_{\textnormal{train}}}{|\mathcal{X}|}\wedge 1}}_{\begin{subarray}{c}\textnormal{statistical constraint due to}\\ \textnormal{limited samples in $\mathcal{X}$}\end{subarray}}\Bigg{\}},$ (5)
where $\widetilde{\delta}=\delta(1+\frac{1}{en})\wedge 1$.
In Theorem 1, we can see the power of any universally valid test is bounded by the minimum of three terms. Here we take a closer look at each of the three terms to better understand how this limit on power arises.
  • First term: computational constraint due to budget $B_{\textnormal{train}}$. The first term is low—that is, not much larger than $1$, meaning that the power cannot be much better than random—whenever the ratio $B_{\textnormal{train}}/n$ is small. This is intuitive: this ratio describes the number of times we are able to call $\mathcal{A}$ on training samples of size $n$, and we need this number to be large in order to distinguish between stability (i.e., large perturbations occur with small probability $\delta^{*}_{\epsilon}\leq\delta$) versus instability (i.e., large perturbations occur with probability $>\delta$).
  • Second term: statistical constraint due to limited samples in $\mathcal{Y}$. The second term shows that power is low whenever we have both of the following: the ratio $N_{\ell}/n$ is not too large, and the ratio $B_{\textnormal{train}}/|\mathcal{Y}|$ is close to 0. For the first quantity, a bound on $N_{\ell}/n$ means that we do not have access to ample labeled data—the available data sampled from $P$ provides only $N_{\ell}/n$ many independent data sets of size $n$, so it is not easy to empirically estimate how often instability may occur. For the second ratio, if $B_{\textnormal{train}}/|\mathcal{Y}|\approx 0$ then this means that we do not have sufficient computational budget to allow an exhaustive search over $\mathcal{Y}$ (i.e., we cannot afford to test the behavior of $\mathcal{A}$ on all possible values of $Y$) to compensate for the limited amount of available labeled data.
  • Third term: statistical constraint due to limited samples in $\mathcal{X}$. The interpretation for this term is analogous to the previous term, with the difference that we are given $N_{\ell}+N_{u}$ many draws of $X$ due to the additional unlabeled data set $\mathcal{D}_{u}$.
In particular, the power can be only slightly higher than $\alpha$ (i.e., only slightly better than random), unless all of the three terms in the minimum are large: to have high power, we would need $B_{\textnormal{train}}\gg n$ to enable sufficiently many calls to $\mathcal{A}$, and moreover, to search over $\mathcal{Y}$ space we need either $N_{\ell}\gg n$ or $|\mathcal{Y}|$ to be relatively small, and similarly to search over $\mathcal{X}$ space we need either $N_{\ell}+N_{u}\gg n$ or $|\mathcal{X}|$ to be relatively small. Note the second and third terms are statistical limits, due to limited data (i.e., $N_{\ell}$ and $N_{u}$), but these statistical limits can be overcomed if $|\mathcal{X}|$ and $|\mathcal{Y}|$ are small enough to permit exhaustive search (under the computation constraint $B_{\textnormal{train}}$).
An unlimited budget?
Comparison with existing work.
4 Discussion
In this paper, we provide a unified power upper bound for any universally valid black-box test for testing algorithmic stability under computational constraints. We reveal the computational budget can be the dominant constraint when the budget $B_{\textnormal{train}}$ is small. At the same time, we illustrate that when the data space is finite, if sample size is limited then certifying stability is possible only if the data spaces $\mathcal{X}$ and $\mathcal{Y}$ are small enough to essentially permit an exhaustive search. In this work, we focus on testing $(\epsilon,\delta)$-stability defined in Definition 1; since a wide range of different stability definitions exist in the literature (see, e.g., Bousquet and Elisseeff, (2002)), an interesting open question is whether there is a unified way to treat different notions of stability under the black-box testing framework.
5 Proof of Theorem 1
The theorem can be separated into three claims and we will prove each of these claims separately. Respectively, these three results reflect the bounds on power that are due to a limited number of calls to $\mathcal{A}$ (a computational limit), and to limited $Y$ data and $X$ data (statistical limits).
Recall that when we run the test ${\widehat{T}}$, the calls $r=1,\dots,{\widehat{r}}$ to the algorithm $\mathcal{A}$ are carried out with data set $\mathcal{D}^{(r)}_{\ell}$ and with random seed $\xi^{(r)}$ (which may be generated in any fashion). Fixing any subset $R\subset[0,1]$, define $\mathcal{E}_{R}$ as the event that $\xi^{(r)}\not\in R$ for all $r$ such that $|\mathcal{D}^{(r)}_{\ell}|=n$—that is, for all calls $r$ to the algorithm during which the training set size is equal to $n$. Let $\textnormal{Leb}(R)$ denote the Lebesgue measure of $R$. Define also the function $f(\xi)=\mathbb{P}\left(|\mathcal{A}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\mid\xi\right)$, the conditional probability of an unstable prediction, as a function of the random seed $\xi$ (here the probability is taken with respect to training and test data sampled i.i.d. from $P$). By definition, we have $\mathbb{E}[f(\xi)]=\delta^{*}_{\epsilon}$, where the expected value is taken with respect to $\xi\sim\textnormal{Unif}[0,1]$.
We will need the following lemma:
Lemma 1.
[…]
With this lemma in place, we are ready to bound the power. Fixing any integer $M\geq 1$, standard measure-theoretic arguments (e.g., Lemma 2.5 of Bennett and Sharpley, (1988)) allow us to construct a partition $[0,1]=S_{1}\cup\dots\cup S_{M}$, where $S_{i}$ is measurable and has Lebesgue measure $1/M$, and $\mathbb{E}[f(\xi)\mid\xi\in S_{i}]=\mathbb{E}[f(\xi)]=\delta^{*}_{\epsilon}$, for each $i\in[M]$.
Proof of Lemma 1.
where $\mathcal{A}_{1}(\mathcal{D};\xi)$ is the function given by $\mathcal{A}_{1}(\mathcal{D};\xi)(x)=1+\epsilon+\max_{\sigma\in S_{|\mathcal{D}|}}\mathcal{A}((\mathcal{D}_{\sigma})_{-|\mathcal{D}|};\xi)(x)$. Here $S_{|\mathcal{D}|}$ denotes the set of all permutations on a set of size $|\mathcal{D}|$, and the notation $(\mathcal{D}_{\sigma})_{-|\mathcal{D}|}$ means that we first permute data set $\mathcal{D}$, i.e., $\left((x_{\sigma(1)},y_{\sigma(1)}),\ldots,(x_{\sigma(|\mathcal{D}|)},y_{\sigma(|\mathcal{D}|)})\right)$, and then remove the last data point. In other words, we have constructed $\mathcal{A}^{\prime}$ so that if the training data set $\mathcal{D}_{n}$ has $n$ data points and $\xi\in R$, then we will inevitably see instability, i.e., we will have $|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(x)-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(x)|>\epsilon$, for any $x$. First we verify that $(\mathcal{A}^{\prime},P,n)$ is not stable: we have
5.2 Statistical limit: bounding power due to limited $Y$ data
For any $y\in\mathcal{Y}$, define $\mathcal{E}_{y}$ as the event that $y\not\in\mathcal{D}^{(1)}_{\ell}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}_{\ell}$, when running the test ${\widehat{T}}$. The key step for this part of the proof is the following lemma:
Lemma 2.
The bound on power is then implied by this lemma, as follows. First, we can assume $|\mathcal{Y}|>B_{\textnormal{train}}$, since otherwise the bound is trivial. Fix any integer $M$ with $B_{\textnormal{train}}<M\leq|\mathcal{Y}|$ (note that $|\mathcal{Y}|$ may be finite or infinite), and let $y_{1},\dots,y_{M}\in\mathcal{Y}$ be distinct. We then have
Proof of Lemma 2.
Fix any $c\in[0,1]$. Define a new distribution
$P^{\prime}=c\cdot(P_{X}\times\delta_{y})+(1-c)\cdot P,$
where $P_{X}$ is the marginal of $X$ under $(X,Y)\sim P$, while $\delta_{y}$ is the point mass at $y$—note that the marginal distribution of $X$ is the same under $P^{\prime}$ as under $P$. Following the construction in the proof of (Kim and Barber,, 2023, Theorem 2), we define also a new algorithm $\mathcal{A}^{\prime}$ as
where the informal notation $y\in\mathcal{D}$ denotes that at least one of the response values in the data points contained in $\mathcal{D}$ is equal to $y$, and where $\mathcal{A}_{1}$ is defined as in the proof of Lemma 1.
5.3 Statistical limit: bounding power due to limited $X$ data
Analogously to our proof for limited $Y$ data, for any $x\in\mathcal{X}$ we define $\mathcal{E}_{x}$ as the event that $x\not\in\mathcal{D}^{(1)}_{\ell}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}_{\ell}$, when running the test ${\widehat{T}}$. For the third bound on power, the key step is to show the following result:
Lemma 3.
With this lemma in place, the proof for the power upper bound due to limited data in $\mathcal{X}$ space is exactly the same as for the previous step (i.e., with $Y$ in place of $X$), so we omit the details and simply turn to the proof of the lemma.
Proof of Lemma 3.
Fix any $c\in[0,1]$. Define a new distribution
$P^{\prime}=c\cdot(\delta_{x}\times P_{Y})+(1-c)\cdot P,$
where the first inequality holds since $((1-c)^{n}-(1-c)^{n+1})\leq\sup_{t\in[0,1]}\{t^{n}-t^{n+1}\}\leq\frac{1}{en}$, and the second inequality holds since $\delta^{*}_{\epsilon}\leq\delta$. Taking any $c>1-\left(\frac{1-\delta(1+\frac{1}{en})}{1-\delta^{*}_{\epsilon}}\right)^{1/n}$, we have $\mathbb{P}_{P^{\prime}}\left\{|\mathcal{A}^{\prime}(\mathcal{D}_{n};\xi)(X_{n+1})-\mathcal{A}^{\prime}(\mathcal{D}_{n-1};\xi)(X_{n+1})|>\epsilon\right\}>\delta$.
where the last step holds by validity of ${\widehat{T}}$ (3), since $(\mathcal{A}^{\prime},P^{\prime},n)$ is not $(\epsilon,\delta)$-stable. We have established that this is true for any $c\in[0,1]$ with $c>1-\left(\frac{1-\delta(1+\frac{1}{en})}{1-\delta^{*}_{\epsilon}}\right)^{1/n}$, this proves the desired bound. ∎
Acknowledgements
Appendix A Formal definition of a black-box test with computational constraints
In this section, we give a formal definition of a black-box test under a computational constraint defined by budget $B_{\textnormal{train}}$. This definition is visualized in Figure 1.
Let $\mathfrak{D}_{\ell}=\cup_{m\geq 0}(\mathcal{X}\times\mathcal{Y})^{m}$ denotes the space of labeled data sets of any size, and let $\mathfrak{D}_{u}=\cup_{m\geq 0}(\mathcal{X})^{m}$ denotes the space of unlabeled data sets of any size.
@@ Appendix A Formal definition of a black-box test with computational constraints
Consider any function ${\widehat{T}}$ that takes as input an algorithm $\mathcal{A}$, a labeled data set $\mathcal{D}_{\ell}\in\mathfrak{D}_{\ell}$, and an unlabeled data set $\mathcal{D}_{u}\in\mathfrak{D}_{u}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})\in\{0,1\}$. We say that ${\widehat{T}}$ is a black-box test with computational budget $B_{\textnormal{train}}>0$ for model training if it can be defined as follows: for some functions $g^{(1)},g^{(2)},\dots$ and $g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
@@
    ${\widehat{T}}=g\big{(}\mathcal{D}_{\ell},\mathcal{D}_{u},(\mathcal{D}_{\ell}^{(r)})_{1\leq r\leq{\widehat{r}}},({\widehat{f}}^{(r)})_{1\leq r\leq{\widehat{r}}},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big{)}.$

### v2

@@ 1 Introduction
That is, $\mathcal{A}$ takes as input $\mathcal{D}=\{(X_{i},Y_{i})\}_{i\in[n]}$, a data set of any size $n$, and returns a fitted model ${\widehat{f}}_{n}:\mathcal{X}\rightarrow\mathbb{R}$, which maps a feature value $x$ to a fitted value, or prediction, ${\widehat{y}}={\widehat{f}}_{n}(x)$.  [^1] Throughout this paper, we implicitly assume measurability for all functions we define, as appropriate. For the algorithm $\mathcal{A}$, we assume measurability of the map $\left((x_{1},y_{1}),\ldots,(x_{n},y_{n}),x\right)\mapsto[\mathcal{A}(\{(x_{i},y_{i})\}_{i\in[n]})](x)$. Algorithmic stability is a notion that quantifies the sensitivity of an algorithm to small changes in the training data $\mathcal{D}$. Since its introduction (Devroye and Wagner,, 1979; Rogers and Wagner,, 1978; Kearns and Ron,, 1997), algorithmic stability has been a central notion in machine learning and statistics. Intuitively, it is a desirable property for learning algorithms—but it also leads to many downstream implications, as well. For instance, it is known that stable algorithms generalize (Kearns and Ron,, 1997; Bousquet and Elisseeff,, 2002; Poggio et al.,, 2004; Shalev-Shwartz et al.,, 2010). It also plays a critical role in a number of different problems, such as model selection (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013; Ren et al.,, 2023), reproducibility (Yu,, 2013), predictive inference (Steinberger and Leeb,, 2023; Barber et al.,, 2021; Liang and Barber,, 2023), selective inference (Zrnic and Jordan,, 2023) and inference for cross-validation error (Austern and Zhou,, 2020; Bayle et al.,, 2020; Kissel and Lei,, 2022).
@@ 1.1 Overview of contributions
In Section 2, we will present some background on algorithmic stability and black-box testing framework under computational constraints. Section 3 presents our main results on the limit of black-box tests for algorithmic stability with computational constraints. A brief discussion is provided in Section 4. We prove the main results in Section 5. Extensions and additional proofs are deferred to the Appendix.
@@ 2 Background and problem setup
The fitted model ${\widehat{f}}_{n}=\mathcal{A}\big{(}\{(X_{i},Y_{i})\}_{i\in[n]};\xi\big{)}$ is now obtained by running algorithm $\mathcal{A}$ on data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, with randomization provided by the argument $\xi\in[0,1]$, which acts as a random seed. Of course, the general definition (2) of a randomized algorithm also accommodates the deterministic (i.e., non-random) setting and we refer any algorithm $\mathcal{A}$ that depends only on the data set $\{(X_{i},Y_{i})\}_{i\in[n]}$, and ignores the second argument $\xi$ as deterministic algorithm. In this paper, we focus on the setting where $\mathcal{A}$ is symmetric in the training data—that is, permuting the order of the $n$ data points does not change the output of $\mathcal{A}$, i.e.,
$\mathcal{A}[(x_{1},y_{1}),\ldots,(x_{n},y_{n});\xi]=\mathcal{A}[(x_{\sigma(1)},y_{\sigma(1)}),\ldots,(x_{\sigma(n)},y_{\sigma(n)});\xi^{\prime}],$
for some coupling of the randomization terms $\xi,\xi^{\prime}\sim\textnormal{Unif}[0,1]$.
@@ 2.1 Algorithmic stability and hypothesis testing formulation
There are many different notions of algorithmic stability in the literature (see, e.g., Bousquet and Elisseeff, (2002) and Appendix A of Shalev-Shwartz et al., (2010) for some comparisons). In this work, we follow Kim and Barber, (2023) and consider testing the following notion of stability:
@@
Although our notation accommodate randomized algorithms, in some cases if we know we are working with deterministic algorithms, we might not need to assume the above strong notion of validity, instead we can require its deterministic conterpart:
$\mathbb{P}_{P}\{{\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=1\}\leq\alpha\textnormal{ for any $(\mathcal{A},P,n)$ that is not $(\epsilon,\delta)$-stable and }\mathcal{A}\textnormal{ is deterministic}.$ (4)
@@ 2.2 Black-box tests with computational constraints
A high-level illustration of a black-box test of interest is provided in Figure 1 and the formal definition is provided in Definition 2. Compared to the definition of black-box test given in Kim and Barber, (2023), here we focus on black-box tests with computational constraints. In particular, the computational budget for training is characterized by $B_{\textnormal{train}}$—the total number of data points that we are allowed to input to the algorithm $\mathcal{A}$, over all calls to $\mathcal{A}$ (for example, when running the test, we can choose to train $\mathcal{A}$ on a data set of size $n$, up to $B_{\textnormal{train}}/n$ many times).
@@
Let $\mathfrak{D}_{\ell}=\cup_{m\geq 0}(\mathcal{X}\times\mathcal{Y})^{m}$ denotes the space of labeled data sets of any size, and let $\mathfrak{D}_{u}=\cup_{m\geq 0}(\mathcal{X})^{m}$ denotes the space of unlabeled data sets of any size. Consider any function ${\widehat{T}}$ that takes as input an algorithm $\mathcal{A}$, a labeled data set $\mathcal{D}_{\ell}\in\mathfrak{D}_{\ell}$, and an unlabeled data set $\mathcal{D}_{u}\in\mathfrak{D}_{u}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})\in\{0,1\}$. We say that ${\widehat{T}}$ is a black-box test with computational budget $B_{\textnormal{train}}>0$ for model training if it can be defined as follows: for some functions $g^{(1)},g^{(2)},\dots$ and $g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
@@
    ${\widehat{T}}(\mathcal{A},\mathcal{D}_{\ell},\mathcal{D}_{u})=g\big{(}\mathcal{D}_{\ell},\mathcal{D}_{u},(\mathcal{D}_{\ell}^{(r)})_{1\leq r\leq{\widehat{r}}},({\widehat{f}}^{(r)})_{1\leq r\leq{\widehat{r}}},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big{)}.$